Find the rule of a linear function f that gives the number of seconds in x weeks.
1
Expert's answer
2013-01-15T08:17:02-0500
1 week = 7 days = 7*24 hours = 7*24*60 minutes = 7*24*60*60 seconds.
To calculate number of seconds in x weeks we should multiply number of seconds in one week by x. So we obtain required rule:
f=7∙24∙60∙60∙x=604800x